# Counting primes with a given primitive root, uniformly

### v1

@@ 1. Introduction
$\sum_{d:\,P^{+}(d)>\log{x}}\frac{\mu(d)}{[\mathbb{Q}(\zeta_{d}):\mathbb{Q}]}=\sum_{d:\,P^{+}(d)>\log{x}}\mu(d)\frac{(d,h)}{d\varphi(d)}+\sum_{\begin{subarray}{c}d:\,P^{+}(d)>\log{x}\\ 2g_{1}\mid d\end{subarray}}\mu(d)\frac{(d,h)}{d\varphi(d)}.$
The first-right hand sum has already been shown to be $O(\frac{\varphi(h)}{h}\frac{\log\log{2|g|}}{\log{x}})$. The second obeys the same bound: In this case,
$\sum_{\begin{subarray}{c}d:\,P^{+}(d)>\log{x}\\ 2g_{1}\mid d\end{subarray}}\mu(d)\frac{(d,h)}{d\varphi(d)}=-\sum_{\ell>\log{x}}\frac{(\ell,h)}{\ell\varphi(\ell)}\sum_{\begin{subarray}{c}d:\,P^{+}(d)<\ell\\ 2g_{1}\mid\ell d\end{subarray}}\mu(d)\frac{(d,h)}{d\varphi(d)}.$ (12)
The right-hand sum on $d$ is empty if $2g_{1}/(2g_{1},\ell)$ has a prime factor at least $\ell$. In all other cases,
@@
But the right-hand side was estimated above as $O(\frac{\varphi(h)}{h}\frac{\log\log{2|g|}}{\log{x}})$.
We conclude that in every case,
$\sum_{d:\,P^{+}(d)>\log{x}}\frac{\mu(d)}{[\mathbb{Q}(\zeta_{d}):\mathbb{Q}]}\ll\frac{\varphi(h)}{h}\frac{\log\log{2|g|}}{\log{x}}.$ (13)
Combining (9), (10), and (13), we arrive at the estimate
$\Pi_{0}(x;g)=A(g)\pi(x)+O\left(\frac{\varphi(h)}{h}\frac{\log\log{2|g|}}{\log x}\pi(x)+x^{3/5}\log|g|\right).$
We will need the error term in “multiplicative form”. Notice that $A_{0}(g)$, as defined in (2), satisfies $A_{0}(g)\asymp\varphi(h)/h$. Recalling the definition (3) of $A_{1}(g)$ in the case when $g_{1}\equiv 1\pmod{4}$, we see that the subtracted term in (3) always has absolute value at most $1$. In fact, that absolute value is at most $1/3$ unless $g_{1}=-3$, in which case $\mu(|g_{1}|)=-1$. Hence, $\frac{2}{3}\leq A_{1}(g)\leq 2$, and
@@
Here in going from the first line to the second, we use that $h/\varphi(h)\ll\log\log{3h}\ll\log\log{x}$ and that $\log|g|\ll x^{1/3}=x^{3/8}/x^{1/24}$. We conclude that
$\Pi_{0}(x;g)=A(g)\pi(x)\left(1+O\left(\frac{\log\log{2|g|}}{\log{x}}\right)\right).$ (14)
@@
Invoking again Lemma 2.1,
@@
Collecting our results,
$E_{1}\ll A(g)\pi(x)\left(\frac{\log\log{2|g|}}{\log{x}}+\frac{\log\log{x}}{\log{x}}\right).$ (16)
We turn now to $E_{2}$. Let $\ell$ be a prime dividing $h$. Then every prime $p\equiv 1\pmod{\ell}$, with $p$ not dividing $g$, satisfies
@@
as $g$ is an $\ell$th power. Hence, in order for a prime $p$ (not dividing $g$) to pass the $\ell$-test, it must be that $p\not\equiv 1\pmod{\ell}$. By assumption, the primes counted in $E_{2}$ pass the $\ell$-test for all $\ell\leq x_{2}$, and hence for all $\ell\leq x^{1/7}$. So if we let $h^{\prime}$ denote the $x^{1/7}$-smooth part of $h$, then each prime $p$ counted in $E_{2}$ has $(p-1,h^{\prime})=1$. Since $p$ also fails the $\ell$-test for some $p\in I_{2}$,
@@
Each prime $p$ counted by the inner sum has the form $p=1+\ell m$. Here $0<m<x/\ell$, and $m$ avoids the residue classes $0\bmod{r}$ for all primes $r\mid h$, $r\leq x^{1/7}$, as well as the residue classes of $-1/\ell\bmod{r}$ for each prime $r<\ell$. Moreover, for each $\ell\in I_{2}$, we have $\ell>x_{2}>x^{1/7}$ as well as $x/\ell\geq x/x_{3}=x_{2}>x^{1/7}$. Applying Brun’s sieve,
@@
We have from (11) that the final product on $r$ is $O(1)$. Thus,
@@
using Mertens’ theorem to estimate the sum on $\ell$. As $A(g)\asymp\varphi(h)/h$, we conclude that
$E_{2}\ll\pi(x)A(g)\left(\frac{\log\log{x}}{\log{x}}+\frac{\log\log{2|g|}}{\log{x}}\right).$ (17)
Finally we consider $E_{3}$. Each $p$ counted in $E_{3}$ has $g^{(p-1)/\ell}\equiv 1\pmod{p}$ for some $\ell>x_{3}$. Thus, the order of $g$ mod $p$ is smaller than $x/x_{3}=x_{2}$, and $p$ divides $g^{m}-1$ for some natural number $m<x_{2}$. The number of distinct prime factors of $g^{m}-1$ is $O(m\log|g|)$, and so
@@
$E_{3}\ll\frac{\pi(x)}{\log{x}}=\frac{\varphi(h)}{h}\pi(x)\cdot\frac{h/\varphi(h)}{\log{x}}\ll A(g)\pi(x)\cdot\frac{\log\log{x}}{\log{x}}.$ (18)
Combining (14), (15), (16), (17), and (18),
$\Pi(x;g)=A(g)\pi(x)\left(1+O\left(\frac{\log\log{x}}{\log{x}}+\frac{\log\log{2|g|}}{\log{x}}\right)\right);$
@@
Now we turn to the proof of Theorem 1.2. We may assume that $|g|$ is sufficiently large. Let $x=\log^{B}|g|$ with $B=19$, and put $W=\prod_{2<p\leq\log x}p$. Denote by $\mathcal{S}$ the set of primes $p\leq x$ with $(g/p)=-1$ and $\gcd(p-1,W)=1$.
First of all, let us estimate the number of elements in $\mathcal{S}$. By inclusion-exclusion,
#$\#\mathcal{S}=\frac{1}{2}\sum_{\begin{subarray}{c}p\leq x,\,p\nmid g\\ (p-1,W)=1\end{subarray}}\left(1-(g/p)\right)$ $=\frac{1}{2}\sum_{\begin{subarray}{c}p\leq x\\ (p-1,W)=1\end{subarray}}\left(1-(g/p)\right)+O(\omega(g))$
$=\frac{1}{2}\sum_{p\leq x}\left(1-(g/p)\right)\sum_{\begin{subarray}{c}d\mid p-1\\ d\mid W\end{subarray}}\mu(d)+O(\log|g|)$
$=\frac{1}{2}\sum_{d\mid W}\mu(d)\sum_{\begin{subarray}{c}p\leq x\\ p\equiv 1\pmod{*}{d}\end{subarray}}\left(1-(g/p)\right)+O(\log|g|)$
$=\frac{1}{2}\sum_{d\mid W}\mu(d)\pi(x;d,1)-\frac{1}{2}\sum_{d\mid W}\mu(d)\sum_{\begin{subarray}{c}p\leq x\\ p\equiv 1\pmod{*}{d}\end{subarray}}(g/p)+O(\log|g|),$
where $\pi(x;d,1)$ denotes the number of primes $p\leq x$ with $p\equiv 1\pmod{d}$. To estimate the first sum above, we appeal to [11, Corollary 13.8], the GRH-conditional prime number theorem for primes in arithmetic progressions, to obtain
$\frac{1}{2}\sum_{d\mid W}\mu(d)\pi(x;d,1)=\frac{1}{2}\sum_{d\mid W}\mu(d)\left(\frac{\mathrm{Li}(x)}{\varphi(d)}+O\left(x^{1/2}\log x\right)\right)$ $=\frac{\tilde{A}_{0}(g)}{2}\mathrm{Li}(x)+O\left(2^{\pi(\log x)}x^{1/2}\log x\right)$
$=\frac{\tilde{A}_{0}(g)}{2}\mathrm{Li}(x)+O\left(x^{1/2+o(1)}\right),$
@@
In addition, we can rewrite the second sum above as
$\frac{1}{2}\sum_{d\mid W}\mu(d)\sum_{\chi\pmod{*}{d}}\sum_{\begin{subarray}{c}p\leq x\\ p\nmid g\end{subarray}}\chi(p)(g/p),$
by the orthogonality relations of Dirichlet characters, where the second summation in the triple sum runs over all Dirichlet characters $\chi\pmod{d}$. It follows that
#$\#\mathcal{S}=\frac{\tilde{A}_{0}(g)}{2}\mathrm{Li}(x)-\frac{1}{2}\sum_{d\mid W}\frac{\mu(d)}{\varphi(d)}\sum_{\chi\pmod{*}{d}}\sum_{\begin{subarray}{c}p\leq x\\ p\nmid g\end{subarray}}\chi(p)(g/p)+O\left(x^{1/2+o(1)}\right).$ (19)
To estimate the triple sum in (19), we recall that $\mathbb{Q}(\sqrt{g})=\mathbb{Q}(\sqrt{g_{1}})$, where $g_{1}\neq 1$ is the unique squarefree integer with $g_{1}(\mathbb{Q}^{\times})^{2}=g(\mathbb{Q}^{\times})^{2}$. Let $\Delta$ be the discriminant of $\mathbb{Q}(\sqrt{g_{1}})$. Then $(g/p)=(\Delta/p)$ for all odd primes $p$ not dividing $g$, and for these primes $p$, $\chi(p)(g/p)$ can be viewed as the value at $p$ of a character $\psi_{\chi,g}\pmod{|\Delta|d}$. The character $\psi_{\chi,g}$ is non-principal unless $\chi$ is induced by the primitive character $(\Delta/\cdot)\pmod{|\Delta|}$. For that to occur, one needs $\Delta\mid d$; in that eventuality, to each $d$ there corresponds exactly one character $\chi\pmod{d}$ for which $\psi_{\chi,g}$ is trivial. All of the $d$ appearing above are odd, squarefree, and divide $W$, so in order for $\Delta$ to divide $d$ we need $\Delta$ to be a squarefree divisor of $W$. This forces $\Delta=g_{1}\equiv 1\pmod{4}$ and requires that $g_{1}\mid W$. By [11, Theorem 13.7], the GRH-conditional estimates for character sums over primes, we have
@@
$=\frac{1_{4\mid(g_{1}-1),\,g_{1}\mid W}}{2}\mathrm{Li}(x)\sum_{g_{1}\mid d\mid W}\frac{\mu(d)}{\varphi(d)}+O\left(2^{\pi(\log x)}x^{1/2}\log x\right)$
$=\frac{1_{4\mid(g_{1}-1),\,g_{1}\mid W}}{2}\cdot\frac{\mu(g_{1})}{\varphi(g_{1})}\mathrm{Li}(x)\sum_{d\mid W/g_{1}}\frac{\mu(d)}{\varphi(d)}+O\left(x^{1/2+o(1)}\right)$
$=\frac{1_{4\mid(g_{1}-1),\,g_{1}\mid W}}{2}\cdot\frac{\mu(g_{1})}{\varphi(g_{1})}\mathrm{Li}(x)\prod_{q\mid W/g_{1}}\left(1-\frac{1}{q-1}\right)+O\left(x^{1/2+o(1)}\right)$
@@
Inserting this estimate in (19) yields
#$\#\mathcal{S}=\frac{\tilde{A}_{0}(g)\tilde{A}_{1}(g)}{2}\mathrm{Li}(x)+O\left(x^{1/2+o(1)}\right).$ (20)
@@
$=\left(1+O\left(\frac{1}{\log\log x}\right)\right)\frac{2C_{2}e^{-\gamma}}{\log\log x}$
and that
@@
is the twin prime constant. Thus, the main term in (20) is of order $\mathrm{Li}(x)/\log\log x$.
@@
[…]
We start with $J_{1}$. Suppose first that $\ell\nmid h$. Applying Lemma 2.1 as in the proof of Theorem 1.1, we see that the count of $p\in\mathcal{S}$ that fail the $\ell$-test for some $\ell\in J_{1}$ is
@@
which is negligible compared to the main term in (20). In the case where $\ell\mid h$, we observe that a prime $p\leq x$ failing the $\ell$-test satisfies $p\equiv 1\pmod{\ell}$ and $\gcd(p-1,W)=1$. For each $\ell\in J_{1}$, the number of such $p\leq x$ is
@@
Since $h\ll\log{|g|}=x^{1/B}$, this is $\ll\mathrm{Li}(x)/(\log\log x)^{2}$, which is also negligible compared to the main term in (20).
Moving on to $J_{2}$, we seek to bound the number of primes $p\in\mathcal{S}$ failing the $\ell$-test for some $\ell\in J_{2}$. Such a prime $p$ certainly satisfies $p\leq x$, $\gcd(p-1,W)=1$, and $p\equiv 1\pmod{\ell}$. Using inclusion-exclusion and invoking [11, Corollary 13.8] again, we find that for each $\ell\in J_{2}$, the number of such $p$ is
@@
$=\tilde{A}_{0}(g)\mathrm{Li}(x)\left(\log\frac{\log y_{2}}{\log y_{1}}+O\left(\frac{1}{\log y_{1}}\right)\right)+O\left(2^{\pi(\log x)}\pi(y_{2})\sqrt{x}\log x\right)$
$=\left(\tilde{A}_{0}(g)\log\frac{B}{B-2}+O\left(\frac{1}{\log\log x}\right)\right)\mathrm{Li}(x)+O\left(x^{1-(1-\log 2+o(1))/\log\log x}\right)$
$=\left(\tilde{A}_{0}(g)\log\frac{B}{B-2}+O\left(\frac{1}{\log\log x}\right)\right)\mathrm{Li}(x),$
where we have made use of the prime number theorem, Mertens’ theorem, and the relation $x=\log^{B}|g|$.
Now we turn to $J_{3}$. As in the treatment of $J_{2}$, we shall only use that a prime $p\in\mathcal{S}$ failing the $\ell$-test satisfies $p\equiv 1\pmod{\ell}$ and that $\gcd(p-1,W)=1$. However, [11, Corollary 13.8] loses its strength in this case, for most $\ell\in J_{3}$ go way beyond $x^{1/2}$. To get around this issue, we resort to the following “arithmetic large sieve” inequality due to Montgomery (see [10, Chapter 3] and [4, §9.4]) to obtain an asymptotically explicit upper bound for the number of primes $p\leq x$ satisfying $p\equiv 1\pmod{\ell}$ and $\gcd(p-1,W)=1$, rather than pursue an asymptotic formula for this count.
@@
Let $Q$ be a positive integer. To each prime $p\leq Q$, associate $\nu(p)<p$ residue classes modulo $p$. For every pair of integers $M,N$, with $N>0$, the number of integers in $[M+1,M+N]$ avoiding the distinguished residue classes mod $p$ for all primes $p\leq Q$ is bounded above by
@@
By the large sieve, the count of $p\leq x$ corresponding to a given $\ell\in J_{3}$ is at most
$\sum_{\begin{subarray}{c}m\leq x/\ell\\ (m,V)=1\\ P^{-}(\ell m+1)>(x/\ell)^{\beta}\end{subarray}}1\leq\left(\frac{x}{\ell}+\left(\frac{x}{\ell}\right)^{2\beta}\right)\left(\sum_{n\leq(x/\ell)^{\beta}}\mu(n)^{2}\prod_{q\mid n}\frac{\nu(q)}{q-\nu(q)}\right)^{-1}$ (21)
@@
$\sum_{n\leq(x/\ell)^{\beta}}\mu(n)^{2}\prod_{q\mid n}\frac{\nu(q)}{q-\nu(q)}\geq\left(\sum_{d\mid V}\mu(d)^{2}\prod_{q\mid d}\frac{2}{q-2}\right)\left(\sum_{\begin{subarray}{c}m\leq(x/\ell)^{\beta}/V\\ (m,V)=1\end{subarray}}\mu(m)^{2}\prod_{q\mid m}\frac{1}{q-1}\right).$ (22)
@@
$\sum_{d\mid V}\mu(d)^{2}\prod_{q\mid d}\frac{2}{q-2}=\prod_{q\mid V}\left(1+\frac{2}{q-2}\right)=\left(1+O\left(\frac{\log\log\log x}{\log\log x}\right)\right)\prod_{q\mid W}\left(1+\frac{2}{q-2}\right).$ (23)

### v2

@@ 2. Notation
$\sum_{d:\,P^{+}(d)>\log{x}}\frac{\mu(d)}{[\mathbb{Q}(\zeta_{d},\sqrt[d]{g}):\mathbb{Q}]}=\sum_{d:\,P^{+}(d)>\log{x}}\mu(d)\frac{(d,h)}{d\varphi(d)}+\sum_{\begin{subarray}{c}d:\,P^{+}(d)>\log{x}\\ 2g_{1}\mid d\end{subarray}}\mu(d)\frac{(d,h)}{d\varphi(d)}.$
The first right-hand sum appeared earlier and was shown to be $O(\frac{\varphi(h)}{h}\frac{\log\log{2|g|}}{\log{x}})$ (see (3.7) and following). To finish off the lemma, it suffices to show that the second right-hand sum is bounded by this same $O$-expression. We rewrite
$\sum_{\begin{subarray}{c}d:\,P^{+}(d)>\log{x}\\ 2g_{1}\mid d\end{subarray}}\mu(d)\frac{(d,h)}{d\varphi(d)}=-\sum_{\ell>\log{x}}\frac{(\ell,h)}{\ell\varphi(\ell)}\sum_{\begin{subarray}{c}d:\,P^{+}(d)<\ell\\ 2g_{1}\mid\ell d\end{subarray}}\mu(d)\frac{(d,h)}{d\varphi(d)}.$ (3.11)
The right-hand sum on $d$ is empty if $2g_{1}/(2g_{1},\ell)$ has a prime factor $p$ at least $\ell$. Indeed, in that case the condition $2g_{1}\mid\ell d$ forces $p\mid d$, contradicting $P^{+}(d)<\ell$. In all other cases, letting $r$ denote a prime number,
@@
To conclude, recall that the right-hand side was estimated as $O(\frac{\varphi(h)}{h}\frac{\log\log{2|g|}}{\log{x}})$ already in (3.10).∎
From (3.2) and (3.3), we have $\Pi_{0}(x;g)=\pi(x)\left(\sum_{d}\mu(d)[\mathbb{Q}(\zeta_{d},\sqrt[d]{g}):\mathbb{Q}]^{-1}\right)+O(x^{3/5}\log|g|)$. Using (3.5) and (3.6) to handle the sum on $d$, we arrive at the estimate
$\Pi_{0}(x;g)=A(g)\pi(x)+O\left(\frac{\varphi(h)}{h}\frac{\log\log{2|g|}}{\log x}\pi(x)+x^{3/5}\log|g|\right).$ (3.12)
Our next lemma puts the error term in “multiplicative form”.
Lemma 3.3.
We have
$\Pi_{0}(x;g)=A(g)\pi(x)\left(1+O\left(\frac{\log\log{2|g|}}{\log{x}}\right)\right).$ (3.13)
Proof.
Notice that $A_{0}(g)$, as defined in (1.2), satisfies $A_{0}(g)\asymp\varphi(h)/h$. Recalling the definition (1.3) of $A_{1}(g)$ in the case when $g_{1}\equiv 1\pmod{4}$, we see that the subtracted term in (1.3) always has absolute value at most $1$. In fact, that absolute value is at most $1/3$ unless $g_{1}=-3$, in which case $\mu(|g_{1}|)=-1$. Hence, $\frac{2}{3}\leq A_{1}(g)\leq 2$, and
@@
Furthermore, $h<x$ (see (3.9)), so that
$\frac{h}{\varphi(h)}\ll\log\log{3h}\ll\log\log{x},$
while (again from (3.9))
$\log|g|\ll x^{1/3}=x^{3/8}/x^{1/24}.$
@@
The assertion (3.13) of Lemma 3.3 now follows from (3.12). ∎
@@
Lemma 3.4.
We have
$E_{1}\ll A(g)\pi(x)\left(\frac{\log\log{2|g|}}{\log{x}}+\frac{\log\log{x}}{\log{x}}\right).$ (3.15)
Proof.
Recall that for a prime $\ell$, we are using $N_{\ell}$ for the number of primes $p\leq x$, $p\equiv 1\pmod{\ell}$, for which $g^{(p-1)/\ell}\equiv 1\pmod{p}$. Invoking Lemma 3.1, and keeping in mind that $\mathbb{Q}(\zeta_{\ell},\sqrt[\ell]{g}):\mathbb{Q}]\gg\ell^{2}/(\ell,h)$ by (3.4), we find that
@@
Collecting our observations yields (3.15).∎
Lemma 3.5.
We have
$E_{2}\ll\pi(x)A(g)\left(\frac{\log\log{x}}{\log{x}}+\frac{\log\log{2|g|}}{\log{x}}\right).$ (3.17)
Proof.
Let $\ell$ be a prime dividing $h$. Then every prime $p\equiv 1\pmod{\ell}$, with $p$ not dividing $g$, satisfies
@@
[…]
@@
Each prime $p$ counted by the inner sum has the form $p=1+\ell m$. Here $0<m<x/\ell$, and $m$ avoids the residue classes $0\bmod{r}$ for all primes $r\mid h$, $r\leq x^{1/7}$, as well as the residue classes of $1/\ell\bmod{r}$ for each prime $r<\ell$. Moreover, for each $\ell\in I_{2}$, we have $\ell>x_{2}>x^{1/7}$ as well as $x/\ell\geq x/x_{3}=x_{2}>x^{1/7}$. Applying Brun’s sieve,
@@
(Here and below, “Brun’s sieve” can be taken to refer to Theorem 2.2 on p. 68 of [8].) We have from (3.8) that the final product on $r$ is $O(1)$. Thus,
@@
using Mertens’ theorem [13, Theorem 2.7(d)] to estimate the sum on $\ell$. Recalling that $A(g)\asymp\varphi(h)/h$, we obtain (3.17). ∎
Lemma 3.6.
We have
$E_{3}\ll A(g)\pi(x)\cdot\frac{\log\log{x}}{\log{x}}.$
Proof.
Each $p$ counted in $E_{3}$ has $g^{(p-1)/\ell}\equiv 1\pmod{p}$ for some $\ell>x_{3}$. Thus, the multiplicative order of $g$ mod $p$ is smaller than $x/x_{3}=x_{2}$, and $p$ divides $g^{m}-1$ for some natural number $m<x_{2}$. The number of distinct prime factors of $g^{m}-1$ is $O(m\log|g|)$, and so
@@
$E_{3}\ll\frac{\pi(x)}{\log{x}}=\frac{\varphi(h)}{h}\pi(x)\cdot\frac{h/\varphi(h)}{\log{x}}\ll A(g)\pi(x)\cdot\frac{\log\log{x}}{\log{x}},$ (3.18)
as desired.∎
Combining (3.13), (3.14), (3.15), (3.17), and (3.18),
$\Pi(x;g)=A(g)\pi(x)\left(1+O\left(\frac{\log\log{x}}{\log{x}}+\frac{\log\log{2|g|}}{\log{x}}\right)\right).$
@@
Now we turn to the proof of Theorem 1.2. We may assume that $|g|$ is sufficiently large. Let $x=\log^{B}|g|$ with $B=19$, and put $W=\prod_{2<p\leq\log x}p$. Denote by $\mathcal{S}$ the set of primes $p\leq x$ with (g/p)=1(g/p)=-1 and $\gcd(p-1,W)=1$.
First of all, let us estimate the number of elements in $\mathcal{S}$. We observe that
$\mathcal{S}=\frac{1}{2}\sum_{\begin{subarray}{c}p\leq x,\,p\nmid g\\ (p-1,W)=1\end{subarray}}\left(1-(g/p)\right)=\frac{1}{2}\sum_{\begin{subarray}{c}p\leq x\\ (p-1,W)=1\end{subarray}}\left(1-(g/p)\right)+O(\omega(g)).$
By inclusion-exclusion, we have
$\sum_{\begin{subarray}{c}p\leq x\\ (p-1,W)=1\end{subarray}}\left(1-(g/p)\right)$ $=\sum_{p\leq x}\left(1-(g/p)\right)\sum_{\begin{subarray}{c}d\mid p-1\\ d\mid W\end{subarray}}\mu(d)$
$=\sum_{d\mid W}\mu(d)\sum_{\begin{subarray}{c}p\leq x\\ p\equiv 1\pmod{*}{d}\end{subarray}}\left(1-(g/p)\right)$
$=\sum_{d\mid W}\mu(d)\pi(x;d,1)-\sum_{d\mid W}\mu(d)\sum_{\begin{subarray}{c}p\leq x\\ p\equiv 1\pmod{*}{d}\end{subarray}}(g/p),$
where $\pi(x;d,1)$ denotes the number of primes $p\leq x$ with $p\equiv 1\pmod{d}$. Hence,
$\#\mathcal{S}=\frac{1}{2}\sum_{d\mid W}\mu(d)\pi(x;d,1)-\frac{1}{2}\sum_{d\mid W}\mu(d)\sum_{\begin{subarray}{c}p\leq x\\ p\equiv 1\pmod{*}{d}\end{subarray}}(g/p)+O(\log|g|),$
since $\omega(g)\leq 2\log|g|$. To estimate the first sum above, we appeal to [13, Corollary 13.8], the GRH-conditional prime number theorem for primes in arithmetic progressions, to obtain
$\sum_{d\mid W}\mu(d)\pi(x;d,1)=\sum_{d\mid W}\mu(d)\left(\frac{\mathrm{Li}(x)}{\varphi(d)}+O\left(x^{1/2}\log x\right)\right)$ $=\tilde{A}_{0}(g)\mathrm{Li}(x)+O\left(2^{\pi(\log x)}x^{1/2}\log x\right)$
$=\tilde{A}_{0}(g)\mathrm{Li}(x)+O\left(x^{1/2+o(1)}\right),$
@@
In addition, we can rewrite
$\sum_{\begin{subarray}{c}p\leq x\\ p\equiv 1\pmod{*}{d}\end{subarray}}(g/p)=\frac{1}{\varphi(d)}\sum_{\chi\pmod{*}{d}}\sum_{\begin{subarray}{c}p\leq x\\ p\nmid g\end{subarray}}\chi(p)(g/p),$
by the orthogonality relations of Dirichlet characters, where the outer sum on the right-hand side runs over all Dirichlet characters $\chi\pmod{d}$. It follows that
$\mathcal{S}=\frac{\tilde{A}_{0}(g)}{2}\mathrm{Li}(x)-\frac{1}{2}\sum_{d\mid W}\frac{\mu(d)}{\varphi(d)}\sum_{\chi\pmod{*}{d}}\sum_{\begin{subarray}{c}p\leq x\\ p\nmid g\end{subarray}}\chi(p)(g/p)+O\left(x^{1/2+o(1)}\right).$ (4.1)
To estimate the triple sum in (4.1), we recall that $\mathbb{Q}(\sqrt{g})=\mathbb{Q}(\sqrt{g_{1}})$, where $g_{1}\neq 1$ is the unique squarefree integer with $g_{1}(\mathbb{Q}^{\times})^{2}=g(\mathbb{Q}^{\times})^{2}$. Let $\Delta$ be the discriminant of $\mathbb{Q}(\sqrt{g_{1}})$. Then $(g/p)=(\Delta/p)$ for all odd primes $p$ not dividing $g$. For these primes $p$, $\chi(p)(g/p)$ can be viewed as the value at $p$ of a character $\psi_{\chi,g}\pmod{|\Delta|d}$. The character $\psi_{\chi,g}$ is non-principal unless $\chi$ is induced by the primitive character $(\Delta/\cdot)\pmod{|\Delta|}$. For that to occur, one needs $\Delta\mid d$; in that eventuality, to each $d$ there corresponds exactly one character $\chi\pmod{d}$ for which $\psi_{\chi,g}$ is trivial. All of the $d$ appearing above are odd, squarefree, and divide $W$, so in order for $\Delta$ to divide $d$ we need $\Delta$ to be a squarefree divisor of $W$. This forces $\Delta=g_{1}\equiv 1\pmod{4}$ and requires that $g_{1}\mid W$. By [13, Theorem 13.7], the GRH-conditional estimates for character sums over primes, we have
@@
$=\frac{1_{4\mid(g_{1}-1),\,g_{1}\mid W}}{2}\mathrm{Li}(x)\sum_{g_{1}\mid d,\,d\mid W}\frac{\mu(d)}{\varphi(d)}+O\left(2^{\pi(\log x)}x^{1/2}\log x\right)$
$=\frac{1_{4\mid(g_{1}-1),\,g_{1}\mid W}}{2}\cdot\frac{\mu(g_{1})}{\varphi(g_{1})}\mathrm{Li}(x)\sum_{d\mid(W/g_{1})}\frac{\mu(d)}{\varphi(d)}+O\left(x^{1/2+o(1)}\right)$
$=\frac{1_{4\mid(g_{1}-1),\,g_{1}\mid W}}{2}\cdot\frac{\mu(g_{1})}{\varphi(g_{1})}\mathrm{Li}(x)\prod_{q\mid(W/g_{1})}\left(1-\frac{1}{q-1}\right)+O\left(x^{1/2+o(1)}\right)$
@@
Inserting this estimate in (4.1) yields
$\mathcal{S}=\frac{\tilde{A}_{0}(g)\tilde{A}_{1}(g)}{2}\mathrm{Li}(x)+O\left(x^{1/2+o(1)}\right).$ (4.2)
@@
$=\left(1+O\left(\frac{1}{\log\log x}\right)\right)\frac{2C_{2}\mathrm{e}^{-\gamma}}{\log\log x}$
by Mertens’ theorem [13, Theorem 2.7(e)], where $\gamma=0.577215$... is the Euler–Mascheroni constant, and that
@@
is the twin prime constant. Thus, the main term in (4.2) is of order $\mathrm{Li}(x)/\log\log x$.
@@
We start with $J_{1}$. Suppose first that $\ell\nmid h$. Applying Lemma 3.1 as in the proof of Theorem 1.1, with the asymptotic relation $\pi(x)\sim\mathrm{Li}(x)$ in mind, we see that the count of $p\in\mathcal{S}$ that fail the $\ell$-test for some $\ell\in J_{1}$ with $\ell\nmid h$ is
@@
which is negligible compared to the main term in (4.2). In the case where $\ell\mid h$, we observe that a prime $p\in\mathcal{S}$ failing the $\ell$-test satisfies $p\equiv 1\pmod{\ell}$ and $\gcd(p-1,W)=1$. For each $\ell\in J_{1}$, the number of such $p\leq x$ is
@@
Since $h\ll\log{|g|}=x^{1/B}$, this is $\ll\mathrm{Li}(x)/(\log\log x)^{2}$, which is also negligible compared to the main term in (4.2).
Moving on to $J_{2}$, we seek to bound the number of primes $p\in\mathcal{S}$ failing the $\ell$-test for some $\ell\in J_{2}$. Such a prime $p$ certainly satisfies $p\leq x$, $\gcd(p-1,W)=1$, and $p\equiv 1\pmod{\ell}$. Using inclusion-exclusion and invoking [13, Corollary 13.8] again, we find that for each $\ell\in J_{2}$, the number of such $p$ is
@@
$=\left(\log\frac{\log y_{2}}{\log y_{1}}+O\left(\frac{1}{\log y_{1}}\right)\right)\tilde{A}_{0}(g)\mathrm{Li}(x)+O\left(2^{\pi(\log x)}\pi(y_{2})\sqrt{x}\log x\right)$
$=\left(\log\frac{B}{B-2}+O\left(\frac{1}{\log\log x}\right)\right)\tilde{A}_{0}(g)\mathrm{Li}(x)+O\left(x^{1-(1-\log 2+o(1))/\log\log x}\right)$
$=\left(\log\frac{B}{B-2}+O\left(\frac{1}{\log\log x}\right)\right)\tilde{A}_{0}(g)\mathrm{Li}(x),$
where we have made use of Mertens’ theorem in the first equality and the prime number theorem and the relation $x=\log^{B}|g|$ in the second equality.
Now we turn to $J_{3}$. As in the treatment of $J_{2}$, we shall only use that a prime $p\in\mathcal{S}$ failing the $\ell$-test satisfies $p\equiv 1\pmod{\ell}$ and that $\gcd(p-1,W)=1$. However, [13, Corollary 13.8] loses its strength in this case, for most $\ell\in J_{3}$ go way beyond $x^{1/2}$. To get around this issue, we resort to the following “arithmetic large sieve” inequality due to Montgomery (see [12, Chapter 3] and [5, §9.4]) to obtain an asymptotically explicit upper bound for the number of primes $p\leq x$ satisfying $p\equiv 1\pmod{\ell}$ and $\gcd(p-1,W)=1$, rather than pursue an asymptotic formula for this count.
@@
Let $Q\geq 1$. To each prime $p\leq Q$, associate $\nu(p)<p$ residue classes modulo $p$. For every pair of integers $M,N$, with $N>0$, the number of integers in $[M+1,M+N]$ avoiding the distinguished residue classes mod $p$ for all primes $p\leq Q$ is bounded above by
@@
By the arithmetic large sieve, the count of $p\leq x$ corresponding to a given $\ell\in J_{3}$ is at most
$\sum_{\begin{subarray}{c}m\leq x/\ell\\ (m,V)=1\\ P^{-}(\ell m+1)>(x/\ell)^{\beta}\end{subarray}}1\leq\left(\frac{x}{\ell}+\left(\frac{x}{\ell}\right)^{2\beta}\right)\left(\sum_{n\leq(x/\ell)^{\beta}}\mu(n)^{2}\prod_{q\mid n}\frac{\nu(q)}{q-\nu(q)}\right)^{-1}$ (4.3)
@@
$\sum_{n\leq(x/\ell)^{\beta}}\mu(n)^{2}\prod_{q\mid n}\frac{\nu(q)}{q-\nu(q)}\geq\left(\sum_{d\mid V}\mu(d)^{2}\prod_{q\mid d}\frac{2}{q-2}\right)\left(\sum_{\begin{subarray}{c}m\leq(x/\ell)^{\beta}/V\\ (m,V)=1\end{subarray}}\mu(m)^{2}\prod_{q\mid m}\frac{1}{q-1}\right).$ (4.4)
@@
$\sum_{d\mid V}\mu(d)^{2}\prod_{q\mid d}\frac{2}{q-2}=\prod_{q\mid V}\left(1+\frac{2}{q-2}\right)=\left(1+O\left(\frac{\log\log\log x}{\log\log x}\right)\right)\prod_{q\mid W}\left(1+\frac{2}{q-2}\right).$ (4.5)